OpenStudy (anonymous):

Last year, a large trucking company delivered 5.5 * 10^5 tons of goods with an average value of $23,000 per ton. What was the total value of the goods delivered? Write the answer in scientific notation. A. 1.265 * 10^9 dollars B. 1.265 * 10^10 dollars C. 1.265 * 10^12 dollars D. 1.265 * 10^11 dollars A, B, C, or D? Please explain

OpenStudy (anonymous):

1.265*10^10 dollars

OpenStudy (mayankdevnani):

5.5*10^5 * 23000 = 1.265*10^10

OpenStudy (anonymous):

23,000 * 5.5 * 10^5 = 126500 * 10^5 = 1.265 * 10^10

OpenStudy (mayankdevnani):

ok @zigzagoon2000

OpenStudy (mayankdevnani):

got it

OpenStudy (anonymous):

Yes, thank you guys.

OpenStudy (mayankdevnani):

welcome